# Unification of Exceptional Points and Transmission Peak Degeneracies in a Highly Tunable Magnon-Photon Dimer

### v2

@@ V Discussion and Outlook
Our magnon-photon platform offers practical pathways towards implementing new TPD-based sensors. Specifically, by including the magnon mode, our system allows for direct observation of external magnetic fields through the frequency detuning $\Delta_{f}$ in the ARP configuration, or external electric fields that modify the analog attenuator, adjusting $\kappa_{y}$ and thus dissipation detuning $\Delta_{\kappa}$ in the RP configuration. In this work, we focus on demonstrating the broad tunability of our architecture, but future research can leverage this platform to engineer robust and optimized sensors that likely have less tunability to avoid parameter hypersensivity issues.
Our theoretical framework separately considers the EPs and TPDs in parameter space. We show that the TPDs do not generally occur at the EPs and provide a mathematical framework to differentiate between the location of all EPs, TPDs, and instability transitions. Remarkably, EPs and TPDs can be characterized using the eigenvalues of the dynamical system matrix through $\Delta_{\lambda}$ and $\Delta_{\lambda}^{2}$. Furthermore, we solidify the role of $J$ and $\kappa_{c}$ as control parameters, which can be used to deliberately place TPDs to positions where the eigenbasis remains complete or even orthogonal, reducing the inherent noise enhancement through the diverging Petermann noise factor. Finally, we present new performance metrics for TPDs using $J$ and $\kappa_{c}$ based on their splitting strength and robustness to model errors, facilitating informed optimization of sensor designs.
@@
The framework and methods developed in this study may extend beyond semi-classical dimers. For example, our system naturally lends itself to studying nonlinear phenomena associated with EPs and TPDs, due to the active gain elements that saturate in the unstable regime. The mathematical tools and experimental methods demonstrated here may be generalized to higher-dimensional lattices and more complex resonators and oscillator networks, potentially revealing new spectral phenomena, topological behaviors, higher-dimensional exceptional surfaces, and EPs [38] or even higher-order TPDs. We leave these exciting steps to future work.
VI Methods
VI.1 Device Description
@@
Our setup includes a vector network analyzer (Quantum Machines OPX+ Quantum Controller) and a local oscillator (SignalCore SC5511A) for collecting the data displayed in Fig 1, 3. Our apparatus consists of a custom-built microwave cavity with mechanically tunable frequency and output-port coupling rates and a custom-built oscillator realized from a 1 mm diameter YIG sphere, placed on top of a co-planar waveguide (Minicircuits CMA-83LN+). The YIG sphere is coupled to the ground using orthogonal wire bonds. Magnon and photon modes are coupled over a meter-scale distance using flexible SMA wires (Minicircuits 086-11SM+). To achieve variable gain amplification and tunable coupling, we place a fixed-gain amplifier (Minicircuits CMA-83LN+), directly followed by a digital attenuator (Vaunix LDA-5018V), and tune the attenuation digitally using the Vaunix Python API. A digital phase shifter (Vaunix LPS-802) modifies the phase for hopping between the cavity and YIG.
@@
We use three digital switches to probe each mode individually during experimental operation without modifying the apparatus. Two Vaunix LSW-802P4T switches (1-input, 4-output) are used on the drive and readout lines to select which mode is being driven or read out. A 1-input, 2-output (Vaunix LSW-802PDT) switch is included on one of the coupling wires to disable or enable coupling when probing each mode individually. When disabled, the signal is routed directly to a 50-$\Omega$ terminator to ensure impedance matching and reduce reflections. Directional couplers (Minicircuits ZADC-13-73-S+) drive and read out each mode, attaching to the self-feedback loops.
@@
Experimental control of parameters is achieved through a Python API. $f_{y}$ is primarily tuned using an external magnetic field controlled by an electromagnet driven by a digital current source (SRS CS580). We offset the frequency of the YIG to 6 GHz using a fixed neodymium magnet and tune $B_{z}$ using the current source. To modify $\kappa_{y}$ with high resolution, we insert an analog voltage-controlled attenautor (Analog Devices HMC812ALC4 integrated with EV1 eval board) on the YIG self-feedback loop, and modify the voltage using precision voltage source (SRS DC205). A full schematic diagram including all components and wiring is included in the Supplementary Information.
VI.2 Parameter Calibration
@@
During an experimental sweep, all the parameters $f_{c},f_{y},\kappa_{c},\kappa_{y},J,\phi$ are necessary for data analysis. Parameters $f_{c},f_{y},\kappa_{c},\kappa_{y}$ are measured during each step of the independent variable, using the RF-switch-based probing method described in Sec. VI.1. We use the experimental apparatus to set $f_{c}=f_{y}$ and $\kappa_{c}=\kappa_{y}$, then couple the two modes together and change the digital phase shifter until the hybridized peaks are the same height in transmission, which defines the phase origin, $\phi=0$. Finally, $J$ is a fitted parameter, extracted from fitting $T_{\text{Coupled}}$ (see Supplementary Information) to each trace of the hybridized data.
VI.3 Experimental Nonidealities
@@
Although the experimental platform enables precise control over key parameters, residual experimental uncertainties persist. For instance, we have observed that modifications of phases can result in changes to effective loss rates due to interference, effectively creating crosstalk between experimental parameters for each mode. However, since we can separately infer $f_{c},f_{y},\kappa_{c},\kappa_{y}$, we can calibrate away any crosstalk between these variables in our system. However, experimental nonidealities cannot be ignored for the RP-TPD experimental sweep, where the voltage-controlled attenuator described in Sec. VI.1 is swept. The voltage-controlled attenuator imparts a significant phase shift while modifying attenuation. This parasitic phase shift causes the YIG frequency $f_{y}$ to drastically change during the RPEP experimental sweep, deviating from the required $\Delta_{f}=0$ path needed to see the RP-TPD. To account for this, we implement a simple digital proportional feedback control system with the current-controlled electromagnet to correct $f_{y}$ and keep $\Delta_{f}=0$ while changing $\Delta_{\kappa}$. The effect of this control system is included in the Supplementary Information. Additionally, for the RP configuration, the maximum $\kappa_{y}$ is limited to $\sim 5\text{ MHz}$ before the YIG resonance is indistinguishable from the system noise floor. Thus, this limits the choices of $J$ and $\kappa_{c}$ for experimental scans to ensure that $\kappa_{y}$ does not breach this maximum threshold before the TPD is reached.
VI.4 Data Analysis
@@
When probing the YIG or cavity individually using the RF-switch-based isolation method, we convert the output data extracted from the OPX+ Quantum Controller into a linear scale and fit the data to a Lorentzian profile. From the fit, we extract $f_{c},f_{y},\kappa_{c},\kappa_{y}$, and $\sigma_{\text{fit}}(f_{c}),\sigma_{\text{fit}}(f_{y}),\sigma_{\text{fit}}(\kappa_{c}),\sigma_{\text{fit}}(\kappa_{y})$ using the elements of the least squares regression covariance matrix. Then, for data where the coupling is enabled, we fit $T_{\text{Coupled}}$ to the data (see Supplementary Information), fitting $J$ and $\sigma_{\text{fit}}(J)$. Then, the equation for $\nu_{\pm}$ is fit to the data using $f_{c},f_{y},\kappa_{c},\kappa_{y},J$ which we report as experimental data in Fig. 3. To quantify the fit uncertainty of $\nu_{\pm}$, we fit $T_{\text{Coupled}}$ to the data $N=10^{4}$ times drawing from a multivariate normal distribution in $f_{c},f_{y},\kappa_{c},\kappa_{y},J$ with uncertainties $\sigma_{\text{fit}}(f_{c}),\sigma_{\text{fit}}(f_{y}),\sigma_{\text{fit}}(\kappa_{c}),\sigma_{\text{fit}}(\kappa_{y}),\sigma_{\text{fit}}(J)$. For each shot in $N$, we calculate $\nu_{\pm}$, yielding a distribution $\mathcal{D}_{\text{fit}}(\nu_{\pm})$ with standard deviation $\sigma_{\text{fit}}(\nu_{\pm})$. The error bars in Fig. 3 are $\sigma_{\text{fit}}(\nu_{\pm})$, a measure of the propagation of uncertainty of the fit parameters to the peak locations of the coupled system.
We also seek to obtain a global uncertainty across the entire experiment, accounting for variations in $J$ and $\kappa_{c}$ across an entire sweep. The average of $J$ and $\kappa_{c}$ across an entire experiment is $\mu_{\text{exp}}(J),\mu_{\text{exp}}(\kappa_{c})$ The experimental standard deviation is $\sigma_{\text{exp}}(J),\sigma_{\text{exp}}(\kappa_{c})$. To represent the set of possible values of $\nu_{\pm}$ under experimental uncertainty in $J,\kappa_{c}$, we fit $\nu_{\pm}$ to the data $N=10^{4}$ times drawing from a multivariate normal distribution in $J$, $\kappa_{c}$ with uncertainties $\sigma_{\text{exp}}(J),\sigma_{\text{exp}}(\kappa_{c})$. For each shot in $N$, we calculate $\nu_{\pm}$, yielding a distribution based on experimental parameters $\mathcal{D}_{\text{exp}}(\nu_{\pm})$ with standard deviation $\sigma_{\text{exp}}(\nu_{\pm})$ Fig. 3(c, f) depicts $\mathcal{D}_{\text{exp}}(\nu_{\pm})$ [purple shaded] for panels Fig. 3iii.
The location of EPs, TPDs, and instability is found using $\mu_{\text{exp}}(J),\mu_{\text{exp}}(\kappa_{c})$. The experimental uncertainty in $J$ and $\kappa_{c}$ also propagates to the location of EPs, TPDs, and instability. Using first order uncertainty propagation, we find $\sigma(\Delta_{f,\kappa}^{\text{EP}})=2\sigma(J)$, $\sigma(\Delta_{f,\kappa}^{\text{TPD}})$ is found using Eq. 24. $\sigma(\Delta_{f}^{\text{NL}})$ is found in the same way, equal to $\sigma(\Delta_{f}^{\text{TPD}})$.
@@ Acknowledgements.
We thank Joe Poissant for his exceptional support in designing and fabricating the 3D microwave cavities. We thank Lorenza Viola, Michiel Burgelman, and Vincent P. Flynn for their contributions to the foundational work upon which this study builds and for valuable discussions about EPs. We thank Hailey A. Mullen and Tian Xia for valuable discussions. We also thank Dr. Mukund Vengalattore for the valuable guidance and support of the project. Startup funds from the Thayer School of Engineering, Dartmouth College, supported this work. We gratefully acknowledge support from the DARPA Young Faculty Award No.  D23AP00192, and from the NSF through Grant DGE-2125733. The views and conclusions in this manuscript are those of the authors. They should not be interpreted as representing the official policies, expressed or implied, of DARPA, NSF, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.
IX Author Contributions
A.S.C., J.S.S.G., and M.F. conceived and designed the experiments, while A.S.C. performed the final set of experiments and calibrations. The theoretical model was conceived by A.S.C. and M.F., and formalized and presented by A.S.C., with S.B. developing and presenting the motivation and derivation of the dynamical matrix. A.S.C. collected, managed, and analyzed data, while generating numerical and analytical simulations, receiving feedback from M.F. All authors jointly validated the results. A.S.C. wrote the manuscript, with copious edits and suggestions from M.F. and edits from J.S.S.G. S.B provided edits and contributed to the introduction, literature review, and supplementary information. M.F. conceived of and supervised the project.
X Competing Interests
@@
XI Additional information
XI.1 Supplementary Information
@@
XI.2 Correspondence and Request for Materials

### v3

@@ IV Conclusion and Outlook
This work presents a highly tunable experimental platform and a unified theoretical framework for systematically investigating EPs and TPDs in coupled oscillator systems. A synthetic gauge field is engineered through the complex coupling term $Je^{i\phi}$, effectively unifying previously distinct two-dimensional configurations of EPs and TPDs into a surface of EPs and TPDs. We report the distinct properties of the EPs and TPDs at $\phi=0,\pi$, and $\pi/2$, with analysis and methods that can be extended to any value of $\phi$. Our magnon-photon platform offers practical pathways towards implementing new TPD-based sensors. Specifically, by including the magnon mode, our system allows for direct observation of external magnetic fields through the frequency detuning $\Delta_{f}$ in the $\phi=\pi$ configuration, or external electric fields that modify the analog attenuator, adjusting $\kappa_{y}$ and thus dissipation detuning $\Delta_{\kappa}$ in the $\phi=0$ configuration. Furthermore, sensors that utilize passive or active feedback systems to remain on the hyperbolic trajectory around the TPDs described for other values of $\phi$ could be used to further enhance sensor response. In this work, we focus on demonstrating the broad tunability of our architecture, but future research can leverage this platform to engineer robust and optimized sensors that likely have less tunability to avoid parameter hypersensitivity issues.
@@
The framework and methods developed in this study may extend beyond semi-classical dimers. For example, our system naturally lends itself to the study of nonlinear phenomena associated with EPs and TPDs, due to the active gain elements that saturate in the unstable regime [35]. The mathematical tools and experimental methods demonstrated here may be generalized to higher-dimensional lattices and more complex resonators and oscillator networks, potentially revealing new spectral phenomena, topological behaviors, higher-dimensional exceptional surfaces, and higher order EPs [20], or even higher-order TPDs. We leave these exciting steps to future work.
V Methods
V.1 Device Description
@@
Our setup includes a vector network analyzer (Quantum Machines OPX+ Quantum Controller) and a local oscillator (SignalCore SC5511A) for collecting the data displayed in Fig. 1, 3, 4. Our apparatus consists of a custom-built microwave cavity with mechanically tunable frequency and output-port coupling rates and a custom-built oscillator realized from a 1 mm diameter YIG sphere, placed on top of a co-planar waveguide (Minicircuits CMA-83LN+). The YIG sphere is coupled to the ground using orthogonal wire bonds. Magnon and photon modes are coupled over a meter-scale distance using flexible SMA wires (Minicircuits 086-11SM+). To achieve variable gain amplification and tunable coupling, we place a fixed-gain amplifier (Minicircuits CMA-83LN+), directly followed by a digital attenuator (Vaunix LDA-5018V), and tune the attenuation digitally using the Vaunix Python API. A digital phase shifter (Vaunix LPS-802) modifies the phase for hopping between the cavity and YIG.
@@
We use three digital switches to probe each mode individually during experimental operation without modifying the apparatus. Two Vaunix LSW-802P4T switches (1-input, 4-output) are used on the drive and readout lines to select which mode is being driven or read out. A 1-input, 2-output (Vaunix LSW-802PDT) switch is included on one of the coupling wires to disable or enable coupling when probing each mode individually. When disabled, the signal is routed directly to a 50-$\Omega$ terminator to ensure impedance matching and reduce reflections. Furthermore, when coupling is diabled, all digital attenuators are set to their maximum attenuation (50 dB) to provide additional isolation. Directional couplers (Minicircuits ZADC-13-73-S+) drive and read out each mode, attaching to the self-feedback loops.
@@
Experimental control of parameters is achieved through a Python API. $f_{y}$ is primarily tuned using an external magnetic field controlled by an electromagnet driven by a digital current source (SRS CS580). We offset the frequency of the YIG to 6 GHz using a fixed neodymium magnet and finely tune $B_{z}$ using the current source. To modify $\kappa_{y}$ with high resolution, we insert an analog voltage-controlled attenautor (Analog Devices HMC812ALC4 integrated with EV1 eval board) on the YIG self-feedback loop, and modify the voltage using precision voltage source (SRS DC205). A full schematic diagram including all components and wiring is included in the Supplementary Information.
V.2 Parameter Calibration
@@
During an experimental sweep, all the parameters $f_{c},f_{y},\kappa_{c},\kappa_{y},J,\phi$ are necessary for data analysis. Parameters $f_{c},f_{y},\kappa_{c},\kappa_{y}$ are measured during each step of the independent variable, using the RF-switch-based probing method described in Sec. V.1. To determine $\phi$, we use the experimental apparatus to set $f_{c}=f_{y}$ and $\kappa_{c}=\kappa_{y}$, then couple the two modes together and change the digital phase shifter until the hybridized peaks are the same height in transmission, which defines the phase origin, $\phi=0$. Finally, $J$ is a fitted parameter, extracted once for each value of loop attenuation and loop phase.
V.3 Experimental Non-idealities
@@
Although the experimental platform enables precise control over key parameters, residual experimental uncertainties persist. For instance, we have observed that modifications of phases can result in changes to effective loss rates due to interference, effectively creating crosstalk between experimental parameters for each mode. However, since we can separately infer $f_{c},f_{y},\kappa_{c},\kappa_{y}$, we can calibrate away any crosstalk between these variables in our system. However, experimental nonidealities cannot be ignored for the RP-TPD experimental sweep, where the voltage-controlled attenuator described in Sec. V.1 is swept. The voltage-controlled attenuator imparts a significant phase shift while modifying attenuation. This parasitic phase shift causes the YIG frequency $f_{y}$ to drastically change during the RPEP experimental sweep, deviating from the required $\Delta_{f}=0$ path needed to see the RP-TPD. To account for this, we implement a simple digital proportional feedback control system with the current-controlled electromagnet to correct $f_{y}$ and keep $\Delta_{f}=0$ while changing $\Delta_{\kappa}$. The effect of this control system is included in the Supplementary Information. Additionally, for the RP configuration, the maximum $\kappa_{y}$ is limited to $\sim 5\text{ MHz}$ before the YIG resonance is indistinguishable from the system noise floor. Thus, this limits the choices of $J$ and $\kappa_{c}$ for experimental scans to ensure that $\kappa_{y}$ does not breach this maximum threshold before the TPD is reached.
V.4 Parameter Control System
Due to the non-idealities described in Sec. V.3, we model the relationship between coil current ($I$), attenuator voltage ($V$), and the measured parameters $(f_{y},\kappa_{y})$ using regression. The pair $(I,V)$ are swept over a dense grid of $N=1200$ points. At each $(I,V)$ we measure $(f_{y},\kappa_{y})$ via a Lorentzian fit of transmission. To capture the weak nonlinearities of the voltage-controlled attenuator we use the feature vector $\bm{\psi}(I,V)=[I,V,IV,V^{2},1,V]^{T}$ and perform least-squares regression using NumPy to obtain a weight matrix $\Theta\in\mathbb{R}^{6\times 2}$ which maps features to predicted outputs as
$\begin{bmatrix}\hat{f}_{y}\\ \hat{\kappa}_{y}\end{bmatrix}=\Theta^{T}\cdot\bm{\psi}(I,V).$ (17)
The model achieves $R^{2}\approx 0.995$ for both $I$ and $V$ and RMSE on the order of $\approx 10\text{ KHz}$.
During an experimental sweep, this model serves as a feedforward map from control inputs $(I,V)$ to predicted outputs $(\hat{f}_{y},\hat{\kappa}_{y})$. The inverse problem is solved via Newton’s method to find the input required to attain a specified target. At each iteration, a new measurement of $(f_{y},\kappa_{y})$ is taken. Every 5 iterations, a new measurement of $(f_{c},\kappa_{c})$ is taken to account for noise in the cavity. We calculate $(\Delta_{f},\Delta_{\kappa})$ and the absolute error relative to the target $(\Delta_{f}^{\text{Target}},\Delta_{\kappa}^{\text{Target}})$ is computed, and a Jacobian is constructed analytically from the regression model. A damped Newton update with gain matrix $K_{p}=\mathrm{diag}([0.8,0.8])$ is applied to $(I,V)$, and the process repeats until the error is within a specified tolerance, $\Delta_{f}^{\text{err}},\Delta_{\kappa}^{\text{err}}=[10^{-5},10^{-5}]$. This protocol enables tracking of arbitrary paths in $(\Delta_{f},\Delta_{\kappa})$ using only a learned model and analytical Jacobian.
V.5 Data Analysis
@@
In the Main Text, we strictly use nondimensionalized parameters to present the data. However, when performing data analysis, we use the dimensional parameters, and scale by $J$ at the end. When probing the YIG or cavity individually using the RF-switch-based isolation method, we convert the output data extracted from the OPX+ Quantum Controller into a linear scale and fit the data to a Lorentzian profile. From the fit, we extract $f_{c},f_{y},\kappa_{c},\kappa_{y}$, and $\sigma(f_{c}),\sigma(f_{y}),\sigma(\kappa_{c}),\sigma(\kappa_{y})$ using the elements of the least squares regression covariance matrix. Then, for data where the coupling is enabled, we fit $\nu_{\pm}$ to the data using $f_{c},f_{y},\kappa_{c},\kappa_{y}$, and $J$, which is determined from Sec. V.2 which we scale by $J$ and report as experimental data in Fig. 3 as $\tilde{\nu}_{\pm}$. To quantify the fit uncertainty of $\nu_{\pm}$, we fit the data $N=10^{4}$ times drawing from a multivariate normal distribution in $f_{c},f_{y},\kappa_{c},\kappa_{y}$ with uncertainties $\sigma(f_{c}),\sigma(f_{y}),\sigma(\kappa_{c}),\sigma(\kappa_{y})$. For each shot in $N$, we calculate $\nu_{\pm}$, yielding a distribution $\mathcal{D}(\nu_{\pm})$ with standard deviation $\sigma(\nu_{\pm})$. The error bars in Fig. 3 are $\sigma(\nu_{\pm})/J$, a measure of the propagation of uncertainty of the fit parameters to the peak locations of the coupled system.
We also seek to obtain a measure of global uncertainty across the entire experiment, accounting for variations in $\kappa_{c}$ across an entire sweep. The average $\kappa_{c}$ is reported as data (in the form $\tilde{\kappa}_{c}$, with experimental uncertainty $\sigma_{\text{exp}}(\tilde{\kappa}_{c})$ equal to the standard deviation of $\kappa_{c}$ across the sweep. The location of EPs, TPDs, and instability transitions are calculated using $J,\kappa_{c}$. $\sigma_{\text{exp}}(\kappa_{c})$ propagates to the location of EPs, TPDs, and instability, which we report in Sec. III. For $\phi=0$ and $\phi=\pi$, a closed form solution for the propagation of uncertainty from $\tilde{\kappa}_{c}$ exists. For other values of $\phi$, which require finding roots of a quartic, a Monte Carlo propagation of uncertainty technique is used instead. The resulting error propagation is relatively small, so we choose to omit it from Fig. 3.
@@ Acknowledgements.
We thank Joe Poissant for his exceptional support in designing and fabricating the 3D microwave cavities. We thank Lorenza Viola, Michiel Burgelman, and Vincent P. Flynn for their contributions to the foundational work upon which this study builds and for valuable discussions about EPs. We thank Hailey A. Mullen, Tian Xia, and Sam Sacerdote for valuable discussions and feedback on the manuscript. We also thank Dr. Mukund Vengalattore for the valuable guidance and support of the project. Startup funds from the Thayer School of Engineering, Dartmouth College, supported this work. We gratefully acknowledge support from the DARPA Young Faculty Award No.  D23AP00192, and from the NSF through Grant DGE-2125733. The views and conclusions in this manuscript are those of the authors. They should not be interpreted as representing the official policies, expressed or implied, of DARPA, NSF, or the US Government. The US Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.
VIII Author Contributions
A.S.C., J.S.S.G., and M.F. conceived and designed the experiments, while A.S.C. performed the final set of experiments and calibrations. The theoretical model was conceived by A.S.C. and M.F., and formalized and presented by A.S.C., with S.B. developing and presenting the motivation and derivation of the dynamical matrix. A.S.C. conceived of the reported TPD figures of merit, as well as non-reciprocal-phase TPDs and the experimental control methods necessary to observe them. A.S.C. collected, managed, and analyzed data, while generating numerical and analytical simulations, receiving feedback from M.F. All authors jointly validated the results. A.S.C. wrote the manuscript, with copious edits and suggestions from M.F. and edits from J.S.S.G. S.B provided edits and contributed to the introduction, literature review, and supplementary information. M.F. conceived of and supervised the project.
IX Competing Interests
@@
X Additional information
X.1 Supplementary Information
@@
X.2 Correspondence and Request for Materials